OpenStudy (anonymous):

Determine if the limit of the sequence (e^x)/(2^n) exists or diverges? Do I have to use the squeeze theorem or should I just evaluate directly?

OpenStudy (zarkon):

I assume you mean \(n\to\infty\)

OpenStudy (anonymous):

Yes the limit as n goes to infinity

OpenStudy (zarkon):

\[e^x\] is just a constant then

OpenStudy (zarkon):

do you know \[\lim_{n\to\infty}\frac{1}{2^n}\]

OpenStudy (anonymous):

0

OpenStudy (zarkon):

yes

OpenStudy (zarkon):

\[\lim_{n\to\infty}\frac{e^x}{2^n}=e^x\lim_{n\to\infty}\frac{1}{2^n}\]

OpenStudy (anonymous):

I don't understand why e^x would be a constant though considering your taking the limit as x tends to infinity. Wouldn't that mean e^x would also tend to infinity because the function e^x approaches infinity for significantly large x?

OpenStudy (zarkon):

\[e^x\] is a constant with respect to n

OpenStudy (anonymous):

Oh I'm sorry I gave you the wrong sequence, yah your totally right though. I meant to type the limit as n tends to infinity of e^n/2^n. I think it diverges.

OpenStudy (zarkon):

that one does diverge

OpenStudy (zarkon):

e>2

OpenStudy (zarkon):

\[\frac{e^n}{2^n}=\left(\frac{e}{2}\right)^n\]

OpenStudy (anonymous):

Okay, that makes sense, I didn't even think to rewrite it! Thank you!

OpenStudy (zarkon):

no problem